# Label-Efficient Grasp Joint Prediction with Point–JEPA

### v1

@@ Abstract
We study whether 3D self-supervised pretraining with a Joint-Embedding Predictive Architecture (Point–JEPA) enables label-efficient grasp joint-angle prediction. Meshes are converted to point clouds and tokenized; a ShapeNet-pretrained Point–JEPA encoder feeds a lightweight multi-hypothesis head trained with winner-takes-all and evaluated by top–logit selection. On DLR–Hand II with strict object-level splits, Point–JEPA improves top–logit RMSE and Coverage@15° in low-label regimes (e.g., 26% lower RMSE at 25% data) and reaches parity at full supervision. This suggests JEPA-style pretraining is a practical lever for data-efficient grasp learning.
@@ II Technical Outline
We convert meshes to point clouds and tokenize them into local patches for transformer processing [13].
We utilize a pretrained JEPA backbone trained on ShapeNet point clouds via Point–JEPA [9]. A block sampler selects spatially contiguous target windows, and a sequencer orders patch centers to keep these windows contiguous in the token sequence, following the JEPA masking/targeting scheme [7]. The context encoder sees masked input; the target encoder sees full input to produce latent targets. A lightweight predictor maps context features to the target space; we regress predictions to stop-gradient targets on masked blocks. The target encoder is an EMA copy of the context encoder; no negatives or input reconstruction are required. After pretraining we discard the predictor and EMA target, retaining the context encoder for contextualized patch features (Fig. 2); attention pooling yields a global object embedding (Fig. 3). For clarity, Fig. 2 depicts single masked patches; training uses block masking.
For grasp prediction, the pooled object embedding is concatenated with the 7D wrist pose and passed to a lightweight head that emits joint-angle hypotheses and logits. Training uses a winner-takes-all / min-over- objective to preserve multi-modality (see Sec. III-D); at inference, we select the top–logit hypothesis [14].
@@ III-A Dataset & split protocol
Data and filtering. We use the DLR–Hand II dataset (non-public; for code/protocol/weights see Footnote Label-Efficient Grasp Joint Prediction with Point–JEPA ) built atop ShapeNet object categories, comprising a range from 10 to 500 unique shapes per category. Each shape provides around 240 grasp samples, each consisting of a 7D wrist pose, a 12D joint configuration, and a grasp quality score. After discarding all grasps with score , our dataset contains on the order of 1,500 meshes and 370,000 total grasps.
Object-level, category-stratified splits. To prevent overfitting and pose leakage from multiple grasps of the same mesh, we split at the object level. Sample-level splits proved overly easy and invalidated the value of SSL-based encoding, and so were not used.
@@
Two split packs and reproducibility. We construct two independent train split packs (A/B) with different manifest seeds; validation fixtures (val, test) are identical across packs. Unless otherwise stated, entries report the mean SD over initialization seeds (0,1) and, where available, over split packs. Due to compute constraints, Pack B was run at anchor budgets (25% and 100%); the low-label budgets (1%, 10%) use Pack A. At the anchor points, Pack A and Pack B yield qualitatively similar trends, so we adopt Pack A as the default and use Pack B to check pack sensitivity.
@@ III-D Joint estimation head & loss
We predict candidate joint configurations for each object embedding and wrist pose, together with a logit vector used to rank hypotheses at test time. Grasping is inherently multi-modal: several distinct joint settings can succeed for the same pose, and single-output MSE collapses these modes to their mean, often yielding an interpolated (and infeasible) grasp (see Fig. 4). To preserve distinct modes, we use a winner-takes-all (WTA) / min-over- objective, a standard approach for ambiguous prediction tasks.
The first term penalizes only the closest hypothesis, avoiding mean-collapse; the cross-entropy term teaches the logits to act as a proxy selector for the winning mode. We then use the same selector at inference (pick the top–logit hypothesis), so evaluation does not rely on an oracle “best-of-”. This mirrors multi-hypothesis practice in trajectory forecasting, where sets of candidates are scored and one is chosen at test time.
In preliminary trials we also considered Mixture Density Networks (MDNs) to model multi-modality probabilistically, but found them brittle in this high-dimensional joint space (training instability and covariance issues) and less aligned with our top-1 selection at inference; the WTA formulation remained simpler and more stable.
All experiments use to balance coverage and compute (brief pilots with were inconclusive; a full sweep is left to future work). Beyond the primary top–logit RMSE, we report two diagnostics that probe the multi-hypothesis mechanism without relying on an oracle at evaluation time: (i) a selection gap (top–logit RMSE minus best-of- RMSE) to assess selector fidelity, and (ii) Coverage@, the fraction of samples for which at least one hypothesis lies within of the ground truth.
@@ III-E Evaluation protocol & metrics
At inference, we choose the head with the largest logit (“top–logit”). The main metric is top–logit RMSE on joint angles. We additionally compute coverage@ (a grasp counts as covered if any head is within the angular threshold) and a diversity proxy (mean pairwise dispersion across heads) to quantify multimodality.
@@ IV Results
All metrics are reported on the fixed test split with object-level stratification, averaged over split packs and seeds where available.
@@ IV-A Main outcomes
Table I summarizes top–logit RMSE (radians; lower is better). Three patterns emerge:
Low labels (1–10%): JEPA pretraining yields consistent gains (– relative error reduction), indicating improved label efficiency in scarce-data regimes.
Moderate labels (25%): The largest improvement appears at 25% ( relative), a “sweet spot” where pretrained context helps most before fully supervised training saturates.
Full labels (100%): Performance is effectively at parity; differences are within seed variance (JEPA scratch), suggesting that with enough supervision, training from random initialization can catch up.
@@ V Conclusion
We investigated whether 3D self-supervised pretraining improves grasp-joint prediction under limited labels. Integrating a Point–JEPA backbone with a winner-takes-all multi-hypothesis head yields consistent gains in low-label regimes on DLR–Hand II, with the largest relative improvement at 25% data. At 100% labels, training from scratch attains parity. Diagnostics indicate that the selector is effective (smaller selection gap) and that Coverage@ improves with pretraining, supporting the inference-aware design.

### v2

@@ Abstract
We study whether 3D self-supervised pretraining with Point–JEPA enables label-efficient grasp joint-angle prediction. Meshes are sampled to point clouds and tokenized; a ShapeNet-pretrained Point–JEPA encoder feeds a multi-hypothesis head trained with winner-takes-all and evaluated by top–logit selection. On a multi-finger hand dataset with strict object-level splits, Point–JEPA improves top–logit RMSE and Coverage@15° in low-label regimes (e.g., 26% lower RMSE at 25% data) and reaches parity at full supervision, suggesting JEPA-style pretraining is a practical lever for data-efficient grasp learning.
@@ II Technical Outline
Meshes are sampled into point clouds and grouped into local patches for transformer processing [13]. Using the Point–JEPA encoder pretrained on ShapeNet [9], we follow the JEPA masking scheme [7]: a block sampler picks spatially contiguous target windows and a simple sequencer orders patch centers so those windows remain contiguous in the token sequence. The context encoder sees masked tokens; the target encoder (an EMA copy) sees the full input and provides stop-gradient latent targets. A lightweight predictor maps context features to the target space. After pretraining we discard the predictor and target encoders and retain the context encoder as the backbone. Attention pooling yields a single object embedding (Fig. 2).
For grasp prediction, the object embedding is concatenated with the 6-DoF hand pose (3D translation + unit quaternion; 7 parameters) and passed to a lightweight head that emits joint-angle hypotheses and logits. Training uses a winner-takes-all / min-over- objective to preserve multi-modality (Sec. III-D); at inference, we select the top–logit hypothesis [14].
@@ III-A Dataset & split protocol
Data and filtering. We use a non-public grasp dataset [15], accessed under educational use, built atop ShapeNet object categories, comprising a range from 10 to 500 unique shapes per category. Each shape provides around 240 grasp samples, each consisting of a 6-DoF hand pose, a 12D joint configuration, and a grasp quality score. Following common practice, we discard grasps with score , yielding on the order of 1,500 meshes and 370,000 total grasps.
Object-level, category-stratified splits. To prevent overfitting and pose leakage from multiple grasps of the same mesh, we split at the object level. Sample-level splits proved overly easy and invalidated the value of SSL-based encoding.
@@
Splits & reproducibility. We use a fixed, object-level split pack (Pack A) for all label budgets. To assess robustness, we built an independent Pack B and verified at 25% and 100% that trends mirror Pack A. For reproducibility, we release code that deterministically regenerates both packs from a local dataset root (see Footnote Label-Efficient Grasp Joint Prediction with Point–JEPA ); we do not publish split JSONs or any third-party file paths/assets.
@@ III-D Joint estimation head & loss
We predict candidate joint configurations for each object embedding and hand pose, together with a logit vector used to rank hypotheses at test time. Grasping is inherently multi-modal: several distinct joint settings can succeed for the same pose, and single-output MSE collapses these modes to their mean, often yielding an infeasible grasp.
To preserve distinct modes, we adopt the standard winner–takes–all (WTA; min-over-) objective—widely used in multimodal regression and structured prediction [16, 17]—and previously applied to dexterous grasping [18]. For completeness, the training objective is:
The first term penalizes only the closest hypothesis (avoiding mean collapse); the cross-entropy term trains the logits to act as a selector for the winning mode. At inference, we use the same selector (pick the top–logit hypothesis), avoiding any oracle “best-of-” evaluation. This matches common multi-hypothesis practice in trajectory forecasting and related settings.
We briefly evaluated Mixture Density Networks (MDNs) as a probabilistic alternative, but observed instability in this high-dimensional joint space and poorer alignment with our top-1 test-time selection; the WTA formulation proved simpler and more stable in practice.
@@ III-E Evaluation protocol & metrics
At inference we select the top–logit hypothesis. The primary metric is top–logit RMSE on joint angles. We also report Coverage@ (a sample is covered if any hypothesis is within of ground truth) and the selection gap (lower is better). Unless stated otherwise, .
@@ IV Results
All results use object-level stratification on the validation split and are averaged over seeds (and split packs where available).
@@ IV-A Main outcomes
Table I reports top–logit RMSE (rad; lower is better).
1–10% labels: consistent gains (8–10% relative), indicating improved label efficiency.
25% labels: largest improvement (26% relative), a “sweet spot” before full supervision saturates.
100% labels: parity within seed variance (pretraining has little impact compared to scratch).
@@ V Conclusion
We investigated whether 3D self-supervised pretraining improves grasp–joint prediction under limited labels. Integrating a Point–JEPA backbone with a winner-takes-all multi-hypothesis head yields consistent gains in low-label regimes on the multi-finger hand dataset, with the largest relative improvement at 25% data. At 100% labels, training from scratch attains parity. Diagnostics indicate that the selector is effective (smaller selection gap) and that Coverage@ improves with pretraining, supporting the inference-aware design.
@@ VI Future work
Acknowledgment
This work is supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence (relAI), sponsored by the Federal Ministry of Research, Technology and Space (BMFTR).